OpenStudy (anonymous):

Determine whether each pair of equations represents perpendicular lines y – x = 3 y = –x + 7

OpenStudy (anonymous):

to check the prduct of the slopes should be -1

OpenStudy (saifoo.khan):

y - x = 3 y + x = 7 ======= 2y = 10 y = 5. Now substitute this value in the equation again to get the value for x.

OpenStudy (anonymous):

how did you get the y= 5 though you add up the 7 and 3 to get 10 but how the 5 come along

OpenStudy (saifoo.khan):

2y = 10 divide 2 from both sides, \[y = \frac{10}{2}\]\[y = 5\]

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

so do i do the same for x

OpenStudy (saifoo.khan):

Now as you have the value for y. now insert in any equation to get the corresponding value of x.

myininaya (myininaya):

Well we are suppose to see if the lines are perpendicular => you need to ask the question are the slopes opposite reciprocals?

OpenStudy (saifoo.khan):

oh man! im blind. Thanks @myininaya for saving the day.

myininaya (myininaya):

=>if they are, we have perpendicular lines =>if they are not, we do not have such lines

myininaya (myininaya):

Well unless we have horizontal and vertical lines going on together

myininaya (myininaya):

But for this case we do not worry about that since we do not

myininaya (myininaya):

Put both of these lines into y=mx+b form The first equation needs to be put in this form The second equation is already in this form

myininaya (myininaya):

m is the slope

OpenStudy (anonymous):

A plot is attached. The lines, y=x+3 in blue and y=-x+7 in red are perpendicular. A plot is not a proof, however, in the case of school book problems a plot is usually correct regarding mutual perpendicularity. Two lines are perpendicular if the coefficient of the x term of one line is equal to the negative of the reciprocal of the x coefficient of the other line. 1=-(1/-1) ?, 1=1 in this case so the lines are perpendicular.